OpenStudy (anonymous):

13.5 Consider the logarithm function \(log=exp^{-1}\) on \(\mathbb{R}\). Prove: a) There are of definition of log is \(\mathbb{R}_{+} :=]0,\infty]\)

OpenStudy (anonymous):

@satellite73 welcome, long time did not seen u :)

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i am not sure what your definition of \(\exp\) is my but my guess is that somewhere in the definition your know that a) \(\exp\) is injective (one to one) so inverse exists and b) the range of \(\exp\) is \(]0,\infty)\) making that the domain of \(\exp^{-1}\)

OpenStudy (anonymous):

@satellite73 pls can you show it mathematically, so that i can give to my mentor, this is my last homework and i need some points for be able to participating in exam.

OpenStudy (anonymous):

i don't know how to show it mathematically other than to say what i wrote above. in general if \(f:A\to B\) is injective, then \(f^{-1}\) exists and the domain of \(f^{-1}\) is the codomain of \(f\)

OpenStudy (anonymous):

ahh you mean thats enough and its solution ?

OpenStudy (anonymous):

you should say that for all \(y\in ]0,\infty)\) exists \(x\in \mathbb{R}\) such that \(\exp(x)=y\) and therefore by definition of the inverse \(\log(y)=x\) but this is a very general fact. if \(f\) has and inverse \(f^{-1}\) it is always the case that the image of \(f\) is the domain of \(f^{-1}\) and the domain of \(f\) is the image of \(f^{-1}\)

OpenStudy (anonymous):

ok satellite thank you very much, its much appreciated

OpenStudy (anonymous):

yw